# SRLM: Human-in-Loop Interactive Social Robot Navigation with Large Language Model and Deep Reinforcement Learning

### v2

@@ IV Methodology
SRLM leverages multiple LLM-based large models and a DRL-based model to provide interactive social robotic services with respect to users’ requirements or feedback. In this framework, the LLM-based large model (LNM) and DRL-based approach (RLNM) are adaptively incorporated with an LLM-based evaluation model (LFM), as shown in Fig 2.
@@ IV-A Human-in-Loop Interactive Mechanism
SRLM drives social robots through a high-level guidance and low-level execution incorporation strategy. Firstly, LLM handles user language input to capture semantic features for global guidance generation. We define three typical social robot tasks (P2P, HG, HF). In the HG task, the social robot navigates to a target point while also maintaining a limited distance with the user until reaching the target location. For the HF task, the robot’s target is updated by the user’s real-time location, and the robot must ensure that the user is within its field of view (FOV) at a comfortable social distance. The basic P2P task involves simply assigning a new target to robot. Moreover, social norm attributes are considered, such as pedestrian-first or robot-first, where the robot will come to a full stop when a pedestrian appears within a fixed distance ($d_{s}$) area under the condition of pedestrian-first. Additionally, user personal preference attributes are collected to select different styles of DRL policy networks in RLNM. Here, we trained three pre-trained policy networks with preferences for large, moderate, and minimal social distance.
Figure 3: An illustration of LNM: The prompt engineering of LNM comprises task description, global guidance, data annotation, initialization, historical data, additional information, and encoded state to directly generate low-level robot actions $[a_{x},a_{y}]$.
@@
Subsequently, global guidances $T_{G}$ are further employed in the following low-level execution blocks: LNM, RLNM, and LFM, where global guidance is described in the prompt engineering of LNM and LFM to supervise LM-based and FB-based action execution. RLNM encodes the target and personal attributes into $\Upsilon$, which can modify the conditional function $\mathcal{C}$ and select the preferred policy network.
@@
Additionally, the LLM block can modify the global guidance or replan new global information based on real-time user feedback or new requests. The human-in-loop interactive mechanism enhances the robustness and flexibility of SRLM, allowing users to adjust robot behaviors based on their feelings during the real-time execution process. For example, if the robot is too closed to pedestrians, users can provide personal feedback to modify the robot’s social distance to a larger value.
IV-B Language Navigation Model
@@
LNM adapts LLM’s supervising ability of context semantic inference to drive a social robot as a low-level motion controller in a human-filled environment. Due to the current requirement of textual information input by LLM, the perception information of the social robot (such as the position of pedestrians) have to be converted into textual information via an image-to-text encoder. Here, the image-to-text encoder translates the robot’s observation state into text descriptions, which mainly include the location and velocity of the robot and observed pedestrians, as well as other features such as orientation and personal radius. However, simply feeding the robot’s observation and action pairs into LLM cannot produce a robust controlling sequence, due to the insensitivity and misunderstanding of LLM regarding a set of numeric values. To enhance the LLM’s inference ability on temporal series, SRLM also implements the prompt engineering $T_{HRI}$, which consists of the following parts: task description $T_{T}$, global guidance $T_{G}$, data annotation $T_{D}$, initialization and historical data $T_{H}$, and additional information $T_{A}$, as shown in Fig. 3.
@@
$\mathrm{T}_{\rm HRI}=f[\mathbf{E}_{\rm S2T}(\mathbf{s}_{\rm t}),LLM(Requests)]$ (1)
@@
IV-C Language Feedback Model
SRLM employs the ability of contextual understanding and inference in LNM to enhance the adaptability of RLNM, aligning and tailoring the pre-trianed DRL policy with different personal preferences and task objectives. Therefore, the integration of DRL-based planner and LM-based planner is facilitated by the LFM, which evaluates both actions and estimates their relative weights as follows:
@@ IV-C Language Feedback Model
The LFM’s GoT queries the action pairs $a^{RL}$ and $a^{LM}$ as input with the prompt engineering of LNM and a new obejective description. Then, $a^{RL}$, $a^{LM}$, and $a^{RL}$ $a^{LM}$ respectively are fed into the next three different vertices via generation edges. Subsequently, $a^{RL}$ and $a^{LM}$ are further evaluated and generated via the first-layer three thoughts to score individual actions as $(s^{i}_{1},s^{i}_{2})$. After the first individually evaluation, the $3^{rd}$-layer vertex is aggregated by two $2^{nd}$-layer thoughts and the $1^{st}$-layer $a^{RL}$ and $a^{LM}$ thoughts to incorporate individual action scores. Finally, the $3^{rd}$-layer vertex generates the relative score thought, which is refined to calculate the combinational scores $(s_{1},s_{2})$.
@@ IV-D Reinforcement Learning Navigation Model
Although LNM adapts the remarkable ability of HRI understanding into navigation decision-making, independent LNM still struggles with uncertainty and the infeasibility of decision-making in complex dynamic environments with continuous space. In contrast, DRL-based RLNM leverages a convergent and efficient policy to address these issues as observed in many works [20, 24, 2]. Therefore, inspired by [2], the DRL-based action execution is employed in SRLM with an ST-graph HRI encoder, LLM-RLHF block, and RLNM block.
@@
$\mathbf{X}_{\mathrm{E}}=\mathcal{F}_{NaviSTAR}(\mathbf{s}_{1},\cdots,\mathbf{s}_{\rm{t}}~{}||~{}\mathcal{G}_{ST})$ (3)
Additionally, the RLHF block is developed to exhibit socially compliant robot behaviors based on [26], encoding human intelligence and supervision in the policy training procedure. For policy training, two different random segments $(\sigma_{1},\sigma_{2})$ are displayed to human supervisors at once. Then, human supervisors must label the segment pair with personal preference $\omega$ as $(\sigma_{1},\sigma_{2},\omega)$ into the data buffer, which is utilized to update the reward neural network $r_{\varphi}$ and the robot policy $\pi$. Lastly, the NaviSTAR [2] planner is developed by RLNM to address interactive social navigation tasks as a Dec-POSMDP paradigm, in which it generates macro-action $\mathbf{u}^{\rm RL}$ and local-action $\mathbf{a}^{\rm RL}$ based on HRI latent embedding $\mathbf{X}_{\mathrm{E}}$ as follows:
@@
To understand human intents and preferences directly from language, SRLM designs an adaptive reward function $r_{LM}$ to enhance the LLM-RLHF training procedure based on [27, 28], which is incorporated with the RLHF reward neural network $r_{\varphi}$ to improve the robustness of the DRL-based action executor. As shown in Fig. 5, the environmental programming specification and personal preference are provided as context prompt engineering to generate the LM reward function corresponding to existing programming formulation and supervisor preference.
Figure 5: LLM-RLHF block: The LLM is introduced in our RLHF training procedure to effectively interpret human supervision from language feedback, wherein the reward functions are generated directly.
Although many DRL-based social navigation approaches [2, 24, 21, 20] demonstrate excellent performance benchmarks for SAN tasks in latent HRI inference and social compliance collision avoidance, the learning algorithms are limited and constrained by training conditions and biased policy patterns. This limitation makes it difficult to maintain sufficient performance in general application scenarios. Hence, a low-level execution decoder is employed to fine-tune the DRL-based planner using the relative weights $(\rm s_{1},s_{2})$ from LFM and the LNM action $\mathbf{a}^{\rm LM}$ as follows:
@@ V-A1 Simulation Setup
We conducted simulation experiments to evaluate the performance of our approach and other ablation models. We developed a human-in-loop interactive social robot navigation environment based on a gym social navigation simulator [2, 24]. In comparison with the original simulator version, user real-time language commands and feedback can be directly implemented to adjust robot behaviors during execution, where both large model blocks (LLM, LNM, and LFM) are configured by GPT-4 [29]. However, the kinematics, dynamics, and other environmental constraints remain the same as in previous works [24].
The default scenario of our experiments involves a robot assisting a user who can talk with robot in real-time, in navigating to a target in an open space among several pedestrians, all simulated by the ORCA policy [19]. There are three main interactive task types in the simulation: human-guiding, human-following, and point-to-point. In the human-guiding task, the social robot guides the user to the target with a physical ribbon connection, necessitating that the robot must maintain a a larger space with other pedestrians. The robot must maintain a suitable distance from the user until the user reaches the destination. The P2P task is the same as in previous simulations, with the target potentially changing halfway through the task based on the user’s real-time command.
All approaches were trained with $1\times 10^{4}$ episodes and tested with 500 random cases for each task, conducted on a desktop with an Intel i9-13900k CPU and an Nvidia 4090 GPU. In each training or testing epoch, a human language requests generator was designed to publish task objectives with personal preferences. Particularly, real-time feedback is stochastically established from the generator with a $50\%$ probability halfway through the current task to update the robot’s goal, user preferences, or other attributes.
@@ V-A2 Baselines and Ablation Models
As shown in Table I, we have set up a comparison of our algorithm with five other baselines or ablation models as follows: (1) A traditional navigation strategy ORCA [19] is utilized as the basic planner, wherein only the global LLM block is maintained to understand high-level human commands for task target establishments; (2) We implement CADRL [30] as the baseline for learning-based approaches, in which the LLM block is also employed for interactive navigation; (3) For the second ablation model, LNM and LFM are detached inside to test the performance of RLNM as SR-RLNM; (4) RLNM and LFM are removed to be viewed as the first ablation model, where the robot is driven only from LNM output as SR-LNM; and (5) For the final one, LFM is replaced by a fixed relative parameter pair $(0.5~{}\&~{}0.5)$ as SR-LFM.
@@ V-A3 Evaluation Metrics
As shown in Fig. 6 and Table I, all methods have been evaluated using $500$ random test cases for each task (total $1500$ cases) individually, with two evaluation metrics: successful rate (SR) and social score (SS) [2]. The SS metric considers various social navigation performance factors such as travel time, collision rate, success rate, discomfort level, and etc.
@@
P2P HG HF AVG P2P HG HF AVG
ORCA [19] $43$ $21$ $20$ 28 $35$ $18$ $24$ 26
CADRL [30] $64$ $47$ $52$ 54 $58$ $52$ $55$ 55
SARL [30] $64$ $47$ $52$ 54 $58$ $52$ $55$ 55
RGL [30] $64$ $47$ $52$ 54 $58$ $52$ $55$ 55
SRNN [20] $64$ $47$ $52$ 54 $58$ $52$ $55$ 55
NaviSTAR [2] $64$ $47$ $52$ 54 $58$ $52$ $55$ 55
SR-RLNM $92$ $78$ $77$ 81 $83$ $67$ $65$ 70
SR-LNM $54$ $69$ $71$ 65 $48$ $52$ $58$ 54
SR-LFM $89$ $82$ $84$ 85 $88$ $74$ $72$ 78
SRLM $\mathbf{94}$ $\mathbf{95}$ $\mathbf{93}$ $\mathbf{94}$ $\mathbf{97}$ $\mathbf{93}$ $\mathbf{95}$ $\mathbf{95}$
V-A4 Quantitative Measurement
Firstly, we analyze the ability of each planner in terms of average trajectory quality with SR and SS metrics. The SR and SS statistic box plots of each planner are shown in Fig. 6. SRLM demonstrates $94\%$ SR and 95 SS ($\mu=95,~{}\sigma=1.85$) higher than others. Then, we conducted an ANOVA tests on SR and SS metrics to examine following hypotheses: ($\mathbf{H_{0}^{SR}}$: $\mu_{1}=\mu_{2}=\cdots=\mu$), where SR null hypothesis $\mathbf{H_{0}^{SR}}$ claims that SR is independent of the selection of planners, and ($\mathbf{H_{1}^{SR}}$: $\exists\mu_{i}\neq\mu;~{}i\in[0,\#\mathcal{A}$]), where the SR alternative hypothesis states that there exists a significant difference among algorithms. Similarly, the hypotheses for SS are defined as follows: (SS null hypothesis $\mathbf{H_{0}^{SS}}$: $\mu_{1}=\mu_{2}=\cdots=\mu$), (SS alternative hypothesis $\mathbf{H_{1}^{SS}}$: $\exists\mu_{i}\neq\mu;~{}i\in[0,\#\mathcal{A}$]).
The linear mixed-effects model is introduced to test the relationship between experimental conditions and objectives in the ANOVA. The ANOVA revealed significant variance among the conditions of planner selection, resulting in a p-value < 0.05 and an F-value of 15410.139 as shown in Table II. Overall, the hypotheses ($\mathbf{H_{1}^{SR}},\mathbf{H_{1}^{SS}}$) are confirmed, while ($\mathbf{H_{0}^{SR}},\mathbf{H_{0}^{SS}}$) are rejected. In other words, the condition of planner selection is a significant factor in navigation performance. Specifically, as shown in Table I and Fig. 6, we observe that SRLM generates the best trajectory for social robot navigation compared to other strategies. The expected means and confidence intervals for all criteria of each planner are summarized by Fig. 6.
TABLE II: : Social Score ANOVA Table
Sum Sq df Mean Sq F-value p-value
Planner 4280451.552 5 856090.310 15410.139 $<0.0001$
Error 499650.014 8994 55.554
Total 4780101.566 8999

### v3

@@ IV Methodology
SALM leverages multiple LLM-based large models and a DRL-based model to provide interactive social robotic services under the dynamically transferable environment that is different with the training configuration of DRL-based LNM robot strategy, with respect to users’ requirements or feedback. In this framework, the LLM-based navigation large model (LNM) and DRL-based approach (RLNM) are adaptively incorporated with an LLM-based evaluation model (LFM), as shown in Fig 2.
Figure 3: An illustration of large language navigation model (LNM): The prompt engineering of LNM comprises task description, global guidance, data annotation, initialization, historical data, additional information, and encoded state to directly generate low-level robot actions $[v_{x},v_{y}]$.
@@ IV-A Human-in-Loop Interactive Mechanism
SALM drives social robots through high-level guidance and low-level execution incorporation strategy. Firstly, LLM handles user language input to capture semantic features for global guidance generation. We define two typical social robot tasks. In the HF task, the robot’s target is updated by the user’s real-time location, and the robot must ensure that the user is within its field of view (FOV) at a comfortable social distance. The basic P2P task involves simply assigning a new target to robot. Moreover, social norm attributes are considered, such as pedestrian-first or robot-first, where the robot will come to a full stop when a pedestrian appears within a fixed distance ($d_{s}$) area under the condition of pedestrian-first.
@@
Subsequently, global guidance $T_{G}$ is further employed in the following low-level execution blocks: LNM, RLNM, and LFM, where global guidance is described in the prompt engineering of LNM and LFM to supervise LM-based and FB-based action execution. RLNM encodes the target and personal attributes into $\Upsilon$, which can modify the conditional function $\mathcal{C}$.
@@
Additionally, the LLM block can modify the global guidance or replan new global information based on real-time user feedback or new requests. The human-in-loop interactive mechanism enhances the robustness and flexibility of SALM, allowing users to adjust robot behaviors based on their feelings during the real-time execution process. For example, if the robot is too closed to pedestrians, users can provide personal feedback to modify the robot’s social distance to a larger value. Moreover, the employment of LNM address the generalization and transfer issues as well, using the robust commonsense inference ability.
IV-B Large Language Navigation Model
@@
LNM adapts LLM’s supervising ability of context semantic inference to drive a social robot as a low-level motion controller in a human-filled environment. Due to the current requirement of textual information input by LLM, the perception information of the social robot (such as the position of pedestrians) has to be converted into textual information via an image-to-text encoder. Here, the image-to-text encoder translates the robot’s observation state into text descriptions, which mainly include the location and velocity of the robot and observed pedestrians, as well as other features such as orientation and personal radius. However, simply feeding the robot’s observation and action pairs into LLM cannot produce a robust controlling sequence due to the insensitivity and misunderstanding of LLM regarding a set of numeric values. To enhance the LLM’s inference ability on temporal series, SALM also implements the prompt engineering $T_{HRI}$, which consists of the following parts: task description $T_{T}$, global guidance $T_{G}$, data annotation $T_{D}$, initialization and historical data $T_{H}$, and additional information $T_{A}$, as shown in Fig. 3.
@@
$\mathrm{T}_{\rm HRI}=\{\mathbf{E}_{\rm S2T}(\mathbf{s}_{\rm t}),~{}LLM(Requests)\}$ (1)
@@
where $\mathbf{E}_{\rm S2T}(\mathbf{s}_{\rm t})$ denotes the textual representation of robot state $\mathbf{s}_{\rm t}$, which is a language description of robot observation information. $LLM(Requests)$ is the high-level guidance from global LLM, representing real-time user commands.
@@
IV-C Large Language Feedback Model
SALM employs the ability of contextual understanding and inference in LNM to enhance the adaptability of RLNM, aligning and tailoring the pre-trianed DRL policy with different personal preferences and task objectives. Therefore, the integration of DRL-based planner and LM-based planner is facilitated by the LFM, which evaluates both actions and estimates their relative weights as follows:
@@ IV-C Large Language Feedback Model
The GoT inference procedure is mainly composed by three distinct interactive inference chains, where the additional inference in the intermediate thoughts are constructed for effective comparison and evaluation. Firstly, the GoT process of LFM verify and evaluate the $\rm a^{RL}$, $\rm a^{LM}$, and the pair of ($\rm a^{RL}$, $\rm a^{LM}$), respectively. The prompt engineering of LNM is fed into LFM as input. Meanwhile, the generation operation of GoT of $1^{st}$-layer is introduced to generate more inference or calculation steps, with a checklist text that describe the verification steps (e.g. check current distance between robot and its destination, or calculate next timestep relative distances among robot and humans). Subsequently, $a^{RL}$ and $a^{LM}$ are further evaluated via the first-layer three thoughts with their reasoning evidences to score individual actions as $(s^{i}_{1},s^{i}_{2})$, comparing the effectiveness of different actions. After $2^{nd}$-layer evaluation, the $3^{rd}$-layer vertex is aggregated by two $2^{nd}$-layer thoughts and the $1^{st}$-layer $a^{RL}$ and $a^{LM}$ thoughts to incorporate individual action scores. Finally, the $3^{rd}$-layer vertex generates the relative score thought, which is refined in the $4^{th}$-layer vertex to calculate the combinational scores $(s_{1},s_{2})$.
@@ IV-D Reinforcement Learning Navigation Model
Although LNM adapts the remarkable ability of HRI understanding into navigation decision-making, independent LNM still struggles with uncertainty and the infeasibility of decision-making in complex dynamic environments with continuous space. In contrast, DRL-based RLNM leverages a convergent and efficient policy to address these issues as observed in many works [29, 31, 6]. Therefore, inspired by [6], the DRL-based action execution is employed in SALM with an ST-graph HRI encoder and RLNM block.
@@
$\mathbf{X}_{\mathrm{E}}=\mathcal{F}_{NaviSTAR}(\mathbf{s}_{1},\cdots,\mathbf{s}_{\rm{t}}~{}||~{}\mathcal{G}_{ST},LLM(Requests))$ (3)
The hybrid spatial-temporal transformer neural network from $\mathcal{F}_{NaviSTAR}$ is employed according to [6] as follows:
$\mathbf{{X}}_{{\rm S}},\mathbf{{X}}_{{\rm T}}=\operatorname{{Trans}_{ST}}(\operatorname{Multi}(\{\mathbf{s}_{1},\cdots,\mathbf{s}_{\rm{t}}\}))$ (4)
$\mathbf{{X}}_{{\rm E}}=\operatorname{{Trans}_{F}}(\operatorname{CMAtten}(\{\mathbf{{X}}_{{S}},\mathbf{{X}}_{{T}}\}))$
where $\rm Multi(\cdot)$ denotes the multi-head attention mechanism within the spatial-temporal transformer network $\rm Trans_{ST}(\cdot)$, which maps the sequential observation data into spatial HRI features $\rm\mathbf{X}_{S}$ and temporal HRI features $\mathbf{X}_{T}$. $\rm CMAtten(\cdot)$ represents the cross-attention block in $\rm Trans_{F}$. The cross-attention block not only fuses above spatial and temporal features but also captures their underlying dependencies, enabling an integrated and context representation as follows:
$\operatorname{Atten}\left(\mathbf{{Q}},\mathbf{{K}},\mathbf{{V}}\right)={\mathrm{softmax}(\frac{\mathbf{{Q}}\mathbf{(}{{\mathbf{K}}})^{\top}}{\sqrt{{d}_{{h}}}})\mathbf{{V}}}$ (5)
$\operatorname{Multi}\left(\mathbf{{Q}},\mathbf{{K}},\mathbf{{V}}\right)=f_{{fc}}({Atten}_{1},\cdots,{Atten}_{{h}})$
$\operatorname{CMAtten}\left(\mathbf{{Q}},\mathbf{{K}},\mathbf{{V}}\right)=\operatorname{Multi}(\mathbf{{Q}}_{{U}},\mathbf{{K}}_{{\bar{U}}},\mathbf{{V}}_{{\bar{U}}})$
where $d_{h}$ is the dimension of attention. $\mathbf{{Q}}_{{U}}$ and $\mathbf{{K}}_{{\bar{U}}}$, $\mathbf{{V}}_{{\bar{U}}}$ present the different modality query or key-value matrix, where $U\in\{\rm\mathbf{X}_{S},~{}\rm\mathbf{X}_{T}\}$ and $\bar{U}$ is the complemental set of $U$.
We modify the NaviSTAR [6] planner as our RLNM to address interactive social navigation tasks within the Dec-POSMDP paradigm, in which it generates macro-action $\mathbf{u}^{\rm RL}$ and local-action $\mathbf{a}^{\rm RL}$ based on HRI latent embedding $\mathbf{X}_{\mathrm{E}}$ as follows:
@@
Although many DRL-based social navigation approaches [6, 31, 28, 29] demonstrate excellent performance benchmarks for SAN tasks in latent HRI inference and social compliance collision avoidance, the learning algorithms are limited and constrained by training conditions and biased policy patterns. This limitation makes it difficult to maintain sufficient performance in general application scenarios. Hence, an execution decoder is employed to incorporate action-pair $(\mathbf{a}^{\rm RL},\mathbf{a}^{\rm LM})$ from both RLNM and LLM, using the relative weights $(\rm s_{1},s_{2})$ from LFM as follows:
@@ V-A1 Simulation Setup
As shown in Fig. 5, we conducted simulation experiments to evaluate the performance of our approach and other ablation models. We developed a human-in-loop interactive social robot navigation environment based on a gym social navigation simulator [6, 8]. In comparison with the original simulator version, user real-time language commands and feedback can be directly implemented to adjust robot behaviors during execution, where both large model blocks (LLM, LNM, and LFM) are configured by GPT-4o [34]. However, the kinematics, dynamics, and other environmental constraints remain the same as in previous works [6].
The default experimental scenario involves a social robot that navigates toward a target in an human-filled open space, while assisting a user to response possible real-time requirements. These human pedestrians each simulated using the ORCA policy [27]. The simulation encompasses two primary interactive task types: human-following (HF) and point-to-point (P2P) navigation tasks. In the HF task, the robot first navigates to meet the user. Once it reaches the user, it continuously updates its destination to remain positioned behind the user, maintaining an appropriate distance until the user arrives at the final destination. For the P2P navigation task, the robot navigates from a starting point to a specified target. Notably, the target may change mid-task based on real-time commands from the user, requiring the robot to adapt its path accordingly.
@@ V-A2 Baselines and Ablation Models
As shown in Table I, we have set up a comparison of our algorithm with five other baselines or ablation models as follows: (1) A traditional navigation strategy ORCA [27] is utilized as the basic planner, wherein only the global LLM block is maintained to understand high-level human commands for task target establishments; (2) We implement CADRL [35] as the baseline for learning-based approaches, in which the LLM block is also employed for interactive navigation; (3) For the second ablation model, LNM and LFM are detached inside to test the performance of RLNM as SA-RLNM; (4) RLNM and LFM are removed to be viewed as the first ablation model, where the robot is driven only from LNM output as SA-LNM; and (5) For the final one, LFM is replaced by a fixed relative parameter pair $(0.5~{}\&~{}0.5)$ as SA-LFM.
@@ V-A3 Evaluation Metrics
As shown in Table I, all methods have been evaluated using $50$ random test cases within 10 humans and 1 user for each task individually, with two evaluation metrics: successful rate (SR) and social score (SS) [6]. The SS metric considers various social navigation performance factors such as travel time, collision rate, success rate, discomfort level, and etc.
@@
P2P HF AVG P2P HF AVG
ORCA [27] $42$ $20$ 31 $35$ $24$ 30
CADRL [35] $64$ $52$ 58 $58$ $55$ 57
SRNN [28] $82$ $68$ 75 $55$ $53$ 54
NaviSTAR [6] $84$ $64$ 74 $70$ $65$ 68
SA-RLNM $84$ $66$ 75 $83$ $65$ 74
SA-LNM $54$ $70$ 62 $48$ $58$ 53
SA-LFM $64$ $62$ 63 $60$ $57$ 59
SALM $\mathbf{90}$ $\mathbf{82}$ $\mathbf{86}$ $\mathbf{85}$ $\mathbf{78}$ $\mathbf{82}$
Figure 5: The illustration of human-in-loop interactive social navigation: The social robot is navigating toward the red star destination with a blue circle user across ten green circle humans. User’s feedback is randomly generated with $50\%$ probability to robot(e.g. change the goal position).
V-A4 Qualitative Analysis
Table I presents a comparative study of our proposed approach alongside five baseline and ablation models, evaluated on two tasks P2P and HF with SR and SS metrics. As a traditional navigation strategy, ORCA utilizes only a global LLM block for high-level human command interpretation. Its relatively low average SR (31) and SS (30) highlight the limitations of classical approaches in complex, socially dynamic environments. For CADRL [35], this learning-based method, is also struggled on performance with an average SR of 58 and SS of 57.
For the ablation model, SA-RLNM, by detaching the LLM-based LNM and the LFM, this variant relies solely on the DRL-based RLNM. It achieves a higher average SR of 75, yet its social score suggests that while the navigation is effective, the overall social compliance might be compromised. For SR-LNM , both the RLNM and LFM are removed, driving the robot solely from the LNM output. The reduced performance (average SR of 62, SS of 53) underscores the necessity of the DRL-based component for robust navigation. The LFM of SR-LFM is replaced with fixed parameters, resulting in an average SR of 63 and SS of 59. This suggests that a dynamic, adaptive evaluation mechanism is crucial for optimal performance compared to a static parameter setting. Finally, SALM is integrating the strengths of both LLM-based and DRL-based approaches with an adaptive evaluation model, SALM achieves the highest average SR (86) and SS (82). In summary, the experimental results indicate that our SALM model significantly outperforms both traditional and learning-based baselines. The ablation studies further validate that each component LLM-based navigation, DRL-based planning, and dynamic language feedback—is vital for achieving high performance in socially-aware navigation tasks.